OpenStudy (anonymous):

-2ab-6c=8d solve for a

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM SOLVE FOR Y 5XY - A = 10Z 5XY - A + (A) = 10Z +( A ) ADD ADDITIVE INVERSE A 1 --- [ 5XY = 10Z + A ] MULTIPLY RECIPROCAL X X ^ THAT WILL CANCEL OUT THE X THERE AND SWITCH IT OVER TO THIS SIDE 1 10Z + A --- [ 5 Y = --------- ] DO LIKEWISE WITH 5 5 X 10Z + A Y = --------- 5X THIS IS SOLVED FOR Y

OpenStudy (anonymous):

-2ab - 6c = 8d -2ac - 6c + 6c = 8d + 6c (add 6c to both sides) -2ac = 8d + 6c -2ac = 8d + 6c ----- ------- (divide both sides by -2c) -2c -2c final answer a = - 8d + 6c -------- 2c

OpenStudy (ivettef365):

-2ab-6c=8d -2ab=8d+6c --- -- -- -2b -2b -2b a=- (8d+6c) / 2b a=-4d-3c/b

OpenStudy (ivettef365):

tabykhan is wrong because he devided by -2c instead of -2b